OpenStudy (anonymous):

Janae distributed a jar of salsa equally among 2 serving bowls. When she finished, each bowl contained 6 ounces of salsa. You can use the equation x over two equals six to represent this problem. Explain what x represents. Im confused about what the x would represent..

OpenStudy (anonymous):

the x = number of ounces in the jar.

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

Word problems are tricky. Read them 2 or 4 times till they click.

OpenStudy (anonymous):

What gives that answer to this word problem is that we know we have a Jar and we know after she poors the salsa in the two bowls there was 6 ounces of salsa in each bowel. the equation is x /2(bowels) = 6. Since we know there is 12 ounce because each bowl got 6 ounce, we know the jar had 12 ounces. If we put 12 where x is we get our answer. \[ \frac{x}{2} = 6 \] \[ \frac{12}{2} = 6 \] So x = number of ounces in the jar.